OpenStudy (anonymous):

Which answer is correct? Question: Solve for d. \[c=\frac{ r(4-d) }{ y }\] ANSWER1: \[d=\frac{ -cy+4r }{ r }\] or mathways answer: ANSWER2: \[d=4-\frac{ cy }{ r }\]

OpenStudy (welshfella):

-4r / -r = 4 and we also have cy / -r = -cy/r

OpenStudy (welshfella):

4 - cy -- r

OpenStudy (anonymous):

I'm slightly confused

OpenStudy (phi):

both answer1 and answer2 are correct. You do not see how to change from one version to the other ?

OpenStudy (anonymous):

do you know which the professor will accept?

OpenStudy (phi):

both are correct. If I were the teacher, I would accept the one you got on your own. If you can't explain it, then I would not accept either. The idea is not to get the answer, the idea is to know how to get the answer.

OpenStudy (phi):

For example, when you look at \[ c=\frac{ r(4-d) }{ y } \] you see the "d" is buried in parens on the right side and it looks messy But you could (should) say, "I can get rid of the y by multiplying both sides by y" that is because you know that y/y is 1 so you would do this \[ c\cdot y = \frac{r(4-d)}{\cancel{y}} \cdot \cancel{y}\]

OpenStudy (phi):

we don't usually bother to show the multiply sign (usually a "dot") if we multiply two "letters" so you have \[ cy = r(4-d) \]

OpenStudy (phi):

you should now think: If I could get rid of the r, that would be progress. if I divide r by r I get 1, and multiplying by 1 can be ignored. But if I divide by r on the right side, I have to do the same on the left side.

OpenStudy (phi):

so you do this \[ \frac{cy}{r} = \frac{r}{r} (4-d) \]

OpenStudy (phi):

on the right side r/r is 1 and 1 times (4-d) is (4-d) so now it is \[ \frac{cy}{r}= 4-d \] we can drop the parens because they are not doing anything.

OpenStudy (phi):

if you add -4 to both sides like this \[ \frac{cy}{r}-4 = 4-4 -d \] and simplify 4-4 to 0, and 0-d is just -d \[ \frac{cy}{r}-4 = -d \] that is almost d all by itself.

OpenStudy (phi):

to change -d to +d, you can multiply it by -1 because -1*-d is d and you have to do the same thing to the other side (which means multiply *everything* on the left by -1 you get \[ -1\cdot \frac{cy}{r} + -1\cdot -4 = d \\ d= -\frac{cy}{r} + 4 \] or, to make it look nicer, we can change the order \[ d = 4 -\frac{cy}{r}\]